OpenStudy (anonymous):

Solve (w-8)^2-80=0 , where w is a real number.

OpenStudy (anonymous):

?

OpenStudy (anonymous):

kk I'm working it out for a sec I'll see if I can help okay @mathgirl2012

OpenStudy (anonymous):

thanks ;)

OpenStudy (anonymous):

http://www.algebra.com/algebra/homework/quadratic/Quadratic_Equations.faq.question.34172.html See if this web page helps if does give medal

OpenStudy (anonymous):

im sorry but that does not help me! :) @The_Jokers_wife

geerky42 (geerky42):

Try to expand \((w-8)^2\) then combine like terms. Then try to factor, or apply quadratic formula.

OpenStudy (anonymous):

aww Okay well I don't know it than I'M SOOOOOO SORRY

OpenStudy (anonymous):

is it \[w=\sqrt{80}+8\] ??

OpenStudy (anonymous):

@geerky42

OpenStudy (cwrw238):

w^2 -16w + 64 -80 = 0

OpenStudy (anonymous):

what??????

OpenStudy (cwrw238):

w^2 - 16w - 16 = 0

OpenStudy (anonymous):

how???????

OpenStudy (cwrw238):

- i did what geerky suggested - expanded the binomial and combined like terms now you have a quadratic to solve

OpenStudy (anonymous):

im still a little confused????

OpenStudy (cwrw238):

(x - 8)^2 = (x - 8)(x - 8) = x^2 - 8x - 8x + 64 = x^2 - 16x + 64

OpenStudy (anonymous):

oooohhhhhh

OpenStudy (anonymous):

so the anwser is ^2 - 16w - 16 = 0

OpenStudy (cwrw238):

- replace x with w

OpenStudy (cwrw238):

no that's not the answer you need to solve for w to solve this equation you need to use the quadratic formula

OpenStudy (anonymous):

\[w^2 - 16w - 16= 0\] \[a= 1, b= -16, c= -16\] \[x=16\frac{ 16\sqrt{256+64} }{ 2 }\] \[x=\frac{ 16\pm \sqrt{320} }{ 2 }\] Is that the anwser @cwrw238

OpenStudy (anonymous):

@cwrw238

OpenStudy (anonymous):

help!!!

OpenStudy (anonymous):

HELP @cwrw238

OpenStudy (anonymous):

PLEASE

OpenStudy (anonymous):

HELP PLEASE

OpenStudy (cwrw238):

take the standard equation a^2 + bx + c = 0 x = [ - b +/- sqrt(b^2 - 4ac)] / 2a

OpenStudy (cwrw238):

compare ur equation with the standard equation; a = 1, b = -16 and c = -16 so replace a , b and c with 1,-16 and -16 in the formula and work it out

OpenStudy (anonymous):

I DID

OpenStudy (cwrw238):

i'll start it off for you - nite the + or - I'll do the + one w = [-(-16) + sqrt( -16^2 - 4 *1*-16)] / 2

OpenStudy (cwrw238):

* i meant note not nite!!

OpenStudy (cwrw238):

you did - good

OpenStudy (anonymous):

can you tell me if I did it right

OpenStudy (cwrw238):

what answers did you get?

OpenStudy (anonymous):

\[x=\frac{ 16\pm \sqrt{320} }{ 2 }\]

OpenStudy (anonymous):

\[x=\frac{ 16\pm 8\sqrt{5} }{ 2 }\]

OpenStudy (anonymous):

??????????????

Nnesha (nnesha):

divide by 2

OpenStudy (cwrw238):

following on w = (16 + sqrt 320) / 2 = (16 + 17.888) / 2 = 16.94

Nnesha (nnesha):

@mathgirl2012 (w-8)^2-80=0 are you doing this one ???

OpenStudy (cwrw238):

the other root is ( 16 - 17.888) / 2

OpenStudy (anonymous):

YES I AM @Nnesha

OpenStudy (anonymous):

8

Nnesha (nnesha):

divide 16 and 8 by 2

OpenStudy (anonymous):

\[\frac{ 2\pm \sqrt{5} }{ 2}\]

OpenStudy (cwrw238):

oh - so they want the answer in surd form?

Nnesha (nnesha):

when you divide by 2 now divide 8 by 2